OpenStudy (anonymous):

how to get the value of a from this equation: a³ - 19a + 30 = 0

OpenStudy (triciaal):

factorization

OpenStudy (anonymous):

can you show me my friend?

Nnesha (nnesha):

have you ever heard p over q method ?

OpenStudy (anonymous):

i didnt

Nnesha (nnesha):

\[\huge\rm \frac{ p }{ q } = \frac{ constant }{ leading ~coefficient }\] what is the constant and leading coefficient in that equation ?

OpenStudy (anonymous):

constant is 30

OpenStudy (anonymous):

how about the leading coefficient?

Nnesha (nnesha):

\[\huge\rm Ax^2 + Bx +C = 0\] coefficient of highest degree variable in this equation A is leading coefficient

OpenStudy (anonymous):

so its 1

OpenStudy (anonymous):

30/1 ? how can we get it?

Nnesha (nnesha):

yep right leading coefficient and constant both number write with plus minus sign \[\huge\rm \frac{ \pm 30 }{ \pm1}\] plus minus 1 would be just one now find all factor of 30

OpenStudy (anonymous):

tell me what can i do to get the value of a immediately please

Nnesha (nnesha):

hmm wait that's not right :(

OpenStudy (anonymous):

i think the value of a is three..

Nnesha (nnesha):

if you need an answer immediately then graph that equation where line cross x-axis's that points would be your answer

Nnesha (nnesha):

highest degree is 3 so there would be 3 zeroes

OpenStudy (anonymous):

i cant im dumb at this thats why i post it D:

Nnesha (nnesha):

can you give me more information about the statement or topic

OpenStudy (anonymous):

this is my geometry :D

Nnesha (nnesha):

aww i'm sorry ;-;